A boy throws a stone straight upward with an initial velocity of 5 m/s from the surface of the Earth. What is the maximum height above the starting point will the stone reach before it starts falling back?

1
Expert's answer
2021-12-07T09:51:45-0500

"h=\frac{v^2-v_0^2}{2g}=\frac{0-5^2}{2\cdot(-9.8)}=1.28\ (m)" . Answer